OpenStudy (anonymous):

Pyramid A is a square pyramid with a base side length of 8 inches and a height of 9 inches. Pyramid B has a volume of 2,880 in3. How many times bigger is the volume of pyramid B than pyramid A? 15 32 54 75 Can you please help me out with this question? @zaibali.qasmi

OpenStudy (anonymous):

as formula for volume of pyramid is V=lwh/3 so here as base is square type so l=w=8in and h=9in so v=576/3=192 now to find the ratio of there sizes just divide 2880 by 192..you will get answer,,that is 15//

OpenStudy (anonymous):

Thank you so much :) @zaibali.qasmi

OpenStudy (anonymous):

my pleasure @shinebrightlikeadimon